OpenStudy (anonymous):

Suppose that the rabbit population on Mr. Jenkins' farm follows the formula: p(t) =(4600t)/(t + 3) where t ≥ 0 is the time (in months) since the beginning of the year. The rabbit population eventually approaches what number?

OpenStudy (anonymous):

4600

OpenStudy (saifoo.khan):

After 1 month, it reaches to: \[p(t) = \frac{4600(1)}{1+3}\] \[p(t) = \frac{4600}{4}\] \[p(t) = 1150.\]

OpenStudy (anonymous):

with your eyeballs. degree of top equals degree of bottom so take ratio of leading coefficients which is 4600

OpenStudy (anonymous):

it tends to 4600

OpenStudy (anonymous):

saifoo! slept that one

OpenStudy (saifoo.khan):

lol, what?

OpenStudy (saifoo.khan):

it says" The rabbit population "eventually" approaches what number? wht does that mean/